### v1

# Repositioning, Ride-matching, and Abandonment in On-demand Ride-hailing Platforms: A Mean Field Game Approach

Abstract.
The on-demand ride-hailing industry has experienced rapid growth, transforming transportation norms worldwide. Despite improvements in efficiency over traditional taxi services, significant challenges remain, including drivers’ strategic repositioning behavior, customer abandonment, and inefficiencies in dispatch algorithms. To address these issues, we introduce a comprehensive mean field game model that systematically analyzes the dynamics of ride-hailing platforms by incorporating driver repositioning across multiple regions, customer abandonment behavior, and platform dispatch algorithms. Using this framework, we identify all possible mean field equilibria as the Karush-Kuhn-Tucker (KKT) points of an associated optimization problem. Our analysis reveals the emergence of multiple equilibria, including the inefficient “Wild Goose Chase” one, characterized by drivers pursuing distant requests, leading to suboptimal system performance. To mitigate these inefficiencies, we propose a novel two-matching-radius nearest-neighbor dispatch algorithm that eliminates undesirable equilibria and ensures a unique mean field equilibrium for multi-region systems. The algorithm dynamically adjusts matching radii based on driver supply rates, optimizing pick-up times and waiting times for drivers while maximizing request completion rates. Numerical experiments and simulation results show that our proposed algorithm reduces customer abandonment, minimizes waiting times for both customers and drivers, and improves overall platform efficiency.
1. Introduction
The on-demand ride-hailing industry has undergone significant expansion in recent years. Uber’s span to over 700 cities worldwide and the pervasive presence of ride-hailing services in China signifies a remarkable shift in transportation norms for people. Although ride-hailing platforms improve efficiency over traditional taxi services, they still encounter challenges and opportunities for enhancement.
One of the primary challenges lies in the complex nature of ride-hailing platforms, which are composed of a large number of agents—primarily drivers—each exhibiting strategic behavior that can significantly impact the system’s overall performance. Specifically, drivers’ repositioning strategies, driven by their anticipation of future demand and competition, play a crucial role in determining the efficiency and equilibrium of the platform. Modeling and analyzing these strategic interactions require sophisticated tools that can capture the dynamic and decentralized decision-making process of drivers.
Another significant challenge is the propensity of customers to abandon the service if they are not matched with a vehicle in a timely manner. This customer abandonment behavior not only affects the user experience but also has implications for the platform’s revenue and market reputation. Addressing this challenge necessitates a deeper understanding of customer preferences and the development of mechanisms that can balance supply and demand more effectively, ensuring that customer needs are met promptly and efficiently.
Finally, the design of efficient dispatch algorithms is paramount to preventing the “Wild Goose Chase” (WGC) equilibrium (Castillo et al., 2023), where drivers chase distant requests inefficiently, leading to increased empty driving time and decreased overall system efficiency. Traditional matching algorithms may induce non-monotonic sojourn times in the driver supply rate, resulting in multiple equilibria and suboptimal outcomes. Innovating in this area requires the development of new algorithms that can dynamically adjust to real-time conditions and optimize the matching process to minimize inefficiencies and maximize service quality.
To model the strategic behavior among a large number of agents (drivers), our study employs the framework of mean field games. This approach simplifies the complex interactions between numerous agents into a representative agent interacting with the mean field, thereby significantly reducing computational complexity. Moreover, when the number of agents is large, this approximation maintains a high degree of accuracy. Given the intricate operational and strategical details involving the platform, passengers, and drivers, constructing an exact model is exceedingly difficult, and precise analysis of such a complex model is nearly impossible. Therefore, we focus on a fundamental mean field game model that captures the core issues. Adopting an agent-based modeling perspective, we consider a closed queueing network comprising a large number of drivers, where each driver’s sojourn time follows a general distribution. The game model we propose is a continuous-time mean field game with finite state-action space where each driver’s decision model is a semi-Markov decision model (semi-MDP). Drawing inspiration from (Kelly, 1989), we apply mean field approximation to express the average sojourn times as functions of the system’s throughput, which in our model corresponds to the driver supply rate—that is, the number of available ride-hailing vehicles per unit time in a specified region. This allows us to transform the problem of finding mean field game equilibria into finding Karush-Kuhn-Tucker (KKT) points for optimization problems. This pivotal idea, initially applied to mean field games in (Courcoubetis and Dimakis, 2023), is further generalized and extended in our work.
Employing the aforementioned mean field game model, this study examines the impact of customer abandonment behavior on platform dispatch algorithms. Specifically, we delve into the classical nearest-neighbor dispatch algorithm, which pairs the closest available vehicle with a customer in need. Our analysis reveals a significant WGC phenomenon. This phenomenon manifests as the existence of multiple equilibria within the system, where “bad equilibrium” results in longer pick-up times compared to “good equilibrium”. In the “bad equilibrium”, distant vehicles are matched with passengers, leading to substantial efficiency losses in the system, which theoretically could become infinitely large. Meanwhile, our analysis reveals that multiple equilibria emerge because the mean pick-up time is a non-monotone function of the driver supply rate. To mitigate the WGC issue, we investigate the effects of introducing matching radius—the maximum distance within which a match is allowed. Our findings indicate that the careful selection of the matching radius is crucial, as an improper choice could still result in multiple equilibria and the persistence of WGC scenarios. The presence of multiple equilibria complicates the task of selecting an optimal matching radius, as there is no explicit solution available. This selection process requires a nuanced consideration of various parameters, particularly the fluctuating number of vehicles in the system.
In light of the difficulties associated with directly determining the optimal matching radius, our study innovatively proposes the Two-Matching-Radius Nearest-Neighbor Dispatch Algorithm, which can be viewed as a generalization of the traditional nearest-neighbor dispatch algorithm. This novel approach designs distinct matching radii for two scenarios: the arrival of new customers and the arrival of new available vehicles. Rather than seeking static optimal matching radii for a fixed system, we solve for optimal adjustment of matching radii, wherein for each given driver supply rate, we find the corresponding optimal matching radii. We derive an explicit relationship between the optimal matching radii and the driver supply rate. Furthermore, we demonstrate that the nearest-neighbor dispatch Algorithm utilizing the optimal adjustment of matching radii results in a sojourn time that is a monotone function of the driver supply rate. This allows us to prove the uniqueness of the mean field equilibrium, thereby theoretically eliminating the WGC equilibrium. By minimizing the sojourn time for each given driver supply rate through our optimal matching radius, our solution not only eradicates undesirable equilibria but also maximizes the system’s throughput, thereby optimizing overall system efficiency for ride-hailing platforms.
In summary, our main contributions are:
  • We introduce a holistic mean field game model to systematically analyze the ride-hailing systems, incorporating key aspects such as driver’s repositioning across multiple regions, customer’s abandonment behavior, and platform’s dispatch algorithms. This unified model allows for the study of their joint effects on system performance.
  • In this general multi-region mean field game framework, we identify all possible mean field equilibria as the Karush-Kuhn-Tucker (KKT) points of an associated optimization problem. Our model reveals the emergence of multiple equilibria, including the inefficient “Wild Goose Chase” one – a well-known matching failure in ride-hailing systems extensively discussed in the literature (Castillo et al., 2023). In this WGC equilibrium, drivers are matched with distant requests and consequently the system efficiency is compromised.
  • We propose a simple yet innovative two-matching-radius nearest-neighbor dispatch algorithm. This algorithm eliminates undesirable equilibria, including the “Wild Goose Chase” one, and ensures a unique mean field equilibrium for the multi-region model.
  • For a single region, our proposed dispatch algorithm minimizes the total pick-up time and waiting time for drivers, thereby maximizing the request completion rate. Numerical experiments and simulation results further demonstrate that the algorithm minimizes customer’s waiting time, reduces customer abandonment, and enhances overall efficiency of the ride-hailing system.
1.1. Literature Review
Our work is related to the following streams of literature.
Repositioning in Ride-hailing Platforms. (Braverman et al., 2019) studied the repositioning of empty cars in a spatial ride-sharing network where a platform controls the repositioning of all vehicles. (Afèche et al., 2023) studied strategical repositioning in ride-hailing networks where drivers are allowed to make repositioning decisions to maximize their earnings. (Benjaafar et al., 2021) and (Gao et al., 2024) considered a combination of decentralized and centralized repositioning to investigate the impact of autonomous vehicles on ride-hailing platforms. Our work is closely related to (Gao et al., 2024) in that we similarly consider a mean field game model and characterize the mean field equilibrium by an optimization problem. The matching procedure are simplified in these literature as they all explicitly or implicitly assume that rider requests are lost if not matched instantly with an available driver. In our model, a detailed matching procedure between riders and drivers is incorporated to the study of the strategical repositioning behavior of drivers.
Ride-matching and Abandonment. Another stream of research close to our work is the study of the performance of matching in ride-hailing systems. (Besbes et al., 2022) considers nearest-neighbor dispatch and uses a spatial queueing model to derive the service rate as a function of pick time and En route time. (Wang et al., 2022) extends the spatial model in (Besbes et al., 2022) to include admission control for matching based on a pickup-time threshold. In addition, abandonment and cancellation on customer’s side are investigated in (Wang et al., 2022). Our model inherits the customer’s abandonment behavior model from (Wang et al., 2022) and similarly consider a two-sided queueing model with open customer-side queue and a closed driver-side queueing network. (Castillo et al., 2023) investigated the “Wild Goose Chase” caused by long pick-up times. (Castillo et al., 2023) suggested that setting a maximum matching radius could improve the system performance. (Yang et al., 2020) studied the impact of matching radius and matching time interval on the system performance. (Wang et al., 2022) and (Yang et al., 2020) both considered batch matching while (Wang et al., 2022) proposed to manipulate the matching radius and matching time interval indirectly through a pickup-time threshold. Different from (Wang et al., 2022) and (Yang et al., 2020), we consider instant matching using nearest-neighbor dispatch, i.e., an arriving customer is immediately matched with the closest idle driver and an arriving driver is immediately matched with the closest waiting customer. Thus, our work views matching radius as the key decision variable of the platform. Our research contribute to this stream of research by introducing a novel two-matching-radius nearest-neighbor dispatch algorithm. The dynamically optimal adjustment of the matching radii proposed in this work offers an solution to the WGC problem.
Mean Field Game. Another area closely related to our work is mean-field games, which investigate the repeated interactions among a large number of strategic agents ((Huang et al., 2006), (Lasry and Lions, 2007)). Our mean-field game model is of finite type, featuring finite state and action spaces. Existing finite mean-field game models in the literature are typically based on either discrete-time Markov Decision Processes (MDPs) (Guo et al., 2023) or continuous-time MDPs with exponentially distributed sojourn times (Neumann, 2020). In our model, we model each driver’s decision-making by continuous-time semi-Markov decision processes (SMDPs) where sojourn times follow general distributions. In fact, our model incorporates a closed queueing network, where drivers can be viewed as “jobs” navigating through the network. Drawing inspiration from (Kelly, 1989), we express sojourn times as functions of driver supply rates. Building on this, we employ the methodology outlined in (Courcoubetis and Dimakis, 2023) to formulate an optimization problem for computing mean-field equilibria. However, the intricate nature of the closed spatial queueing network inherent in ride-hailing systems generally results in sojourn times that are not monotone functions of throughput, thereby violating the monontone assumptions put forth in (Kelly, 1989) and (Courcoubetis and Dimakis, 2023). To address this, we extend their methodologies to accommodate non-monotone scenarios. Moreover, by optimally designing the dispatch algorithm and matching radius, we can ensure that sojourn times become monotone, thereby validating the monotone assumptions of (Kelly, 1989) and (Courcoubetis and Dimakis, 2023).
1.2. Paper Organization
In Section 2, we describe each driver’s continuous-time semi-Markov Decision Process model and formulate a mean field game model. In Section 3, we investigate the ride-matching algorithm under abandonment behavior, derive the sojourn times as exact functions of driver supply rates. In Section 4, optimal matching radius design and a solution to eliminate the “Wild Goose Chase” equilibrium are provided . In Section 5, we conduct simulations to evaluate the performance our proposed dispatch algorithm. We conclude the paper in Section 6.
2. Mean Field Game Model
This section introduces a mean-field game model that focuses on modeling drivers’ strategic repositioning behavior across multiple regions. The details concerning customer behavior and platform strategy will be discussed in subsequent sections.
2.1. A Semi-Markov Decision Process
We consider a closed network comprising a total mass $m$ of drivers and $N$ regions, each region labeled by numbers $1,2,\cdots,N$. These regions can be viewed as distinct geographic areas within a city. The drivers are modeled by a continuous-time semi-Markov decision process (semi-MDP). The state space, denoted by $S=\{1,2,\cdots,N\}$, signifies the driver’s current location. The action space, denoted by $A=\{1,2,\cdots,N\}$ , indicates the driver’s intended destination. Hence, the semi-MDP we consider is a special one with $S=A$. For a given state-action pair $(i,k)\in S\times A=S^{2}$, $k=i$ implies that the driver chooses to provide services within region $i$, while $k\not=i$ indicates that the driver is repositioning to region $k$.
Transition Ptobability. The transition probabilities are encapsulated in the tensor $P=(P^{k}_{ij})_{i,j\in S,k\in A}$, where the subscripts denote the state transition pairs and the superscript indicates the action. $P^{k}_{ij}$ denotes the probability of transitioning from state $i$ to state $j$ under action $k$. Let $q_{ik}$ denote the probability that a passenger in region $i$ has a destination in region $k$. Thus, the matrix $(q_{ik})_{(i,k)\in S\times A}$ is also referred to as the demand matrix. Using the demand matrix, we are able to define the transition matrix of the semi-MDP. The transition probability from state $i$ to state $j$ for any driver with the state-action pair $(i,k)$ is $P_{ij}^{k}=\delta_{jk}$ when $i\not=k$, where $\delta_{jk}=1$ if $j=k$ and 00 otherwise. For the state-action pair $(i,i)$, the transition probability from state $i$ to state $j$ is $P_{ij}^{k}=q_{ik}$. In this context, for a driver in state $i$ taking action $k$ where $k\not=i$, the driver opts to reposition to region $k$, leading to a state transition to region $k$ with probability one. When $k=i$, the driver decides to provide services within the current region, and the state transitions in accordance with the passenger demand matrix. In summary, the transition probabilities $(P^{k}_{ij})_{i,j\in S,k\in A}$ are given by,
$P_{ij}^{k}=\begin{cases}q_{ik}&\text{ if }i=j,\\ \delta_{jk}&\text{ if }i\not=j.\end{cases}$
Sojourn Time. For state-action pair $(i,k)$, it takes a certain amount of time for a driver to complete action $k$ when they are in state/location $i$ (to serve a customer in region $i$ when $k=i$ or reposition to $k$ if $k\not=i$ ). This is also referred to as the sojourn time for the state-action pair $(i,k)$. Let $T_{ik}$ denote the mean time taken for a driver to finish action $k$ from state/location $i$. The sojourn time is a result of interactions across the platform, customers, and drivers, encompassing elements such as the driver’s wait time until a match, the duration of pick-ups, and travel times between locations. Factors influencing the determination of sojourn time include customer behavior, platform matching policies, and drivers’ strategies. Under appropriate assumptions and within specific customer behavior models and platform matching policies, we will show that the mean sojourn time can be calculated based on the distribution of driver masses over the state-action space and their respective mass rates.
Steady State Analysis. We consider a non-atomic game model where a representative driver makes decisions against the mass behavior of all other drivers, characterized by the mass distribution over $S\times A$. This representative driver is negligible and exerts no influence on the system, including the mass distribution and corresponding mass rate. This approach is reasonable given the enormous number of drivers in the system, their anonymity, symmetry, and the sole differentiation based on their locations. We focus on stationary equilibria and study stationary systems, which is sensible since each driver perceives the system to be in a steady state, enabling observation of mass distributions and estimation of sojourn times effectively. Our approach involves a fluid-based analysis that operates with the averages of the sojourn times and associated metrics, a methodology commonly employed in related literature (Braverman et al., 2019; Besbes et al., 2022; Wang et al., 2022; Courcoubetis and Dimakis, 2023).
[…]
A system characterized by mass distribution $\mu$ and corresponding mass rate vector $x$ is considered stationary when $x$ satisfies the following balance equation,
(1) $\sum_{k\in A}x_{ik}=\sum_{{k\in S},k\not=i}x_{ki}+\sum_{k\in S}x_{kk}q_{ki},\>\forall i\in S$
which is interpreted as the mass flow out of state/location $i$ (on the left-hand side) should equal the mass flow into state/location $i$ (on the right-hand side), where the latter is composed of the flow of empty cars for repositioning and the flow of cars carrying passengers. Henceforth, unless explicitly stated otherwise, we focus on stationary systems.
Now, let us proceed to describe how to determine the sojourn times in a stationary system. The mean sojourn time for any $(i,k)\in S\times A$ is firstly dependent on the mean travel time $t_{ik}$ of a requested trip from region $i$ to region $k$ (similarly, $t_{ii}$ is the mean travel time for a requested trip within the region $i$). When $k\not=i$, the vehicle repositions and the mean sojourn time $T_{ik}$ equals the mean travel time $t_{ik}$. When $k=i$, the sojourn time consists of three components: waiting time to get matched to a customer, pick-up time and en route time, with means denoted by $w_{i}^{d}$, $\tau_{i}$ and $t_{i}$ respectively. The mean en route time $t_{i}$ can be computed from the demand matrix and mean travel time matrix $(t_{ij})_{i,j\in S}$ by $t_{i}=\sum_{j\in S}t_{ij}q_{ij}$. The mean waiting time of a driver $w_{i}^{d}$ and the mean pickup $\tau_{i}$ are influenced by both the customers’ arrival process and the platform’s matching policy. We will investigate these factors in the subsequent subsection. In fact, we can conceptualize that, based on the customer behavior model and platform matching policy, we are able to establish in each region relationships between the driver’s average waiting time, pick-up time, and mass rate. More specifically, the following assumption (that will be validated by the analysis) simply postulates that we can derive all other quantities as functions of the mass rate in each region. We use this assumption when we define the underlying mean-field game.
Assumption 1.
For a given vector of customer arrival rates and platform matching policy, when the system is in a stationary state, the driver’s mean waiting time $w_{i}$ and mean pick-up time $\tau_{i}$ are positive and can be written as functions of the mass rate of available drivers $x_{ii}$, for every $i\in S$. Thus, the mean total sojourn time $T_{ii}$ can also be written as a function of $x_{ii}$, i.e.,
$T_{ii}(x_{ii})=w_{i}^{d}(x_{ii})+\tau_{i}(x_{ii})+t_{i},$
where $t_{i}$ is a constant such that $t_{i}=\sum_{j\in S}t_{ij}q_{ij}$. Moreover, we further assume that as $x_{ii}\rightarrow b_{i}$, $w_{i}^{d}(x_{ii})\rightarrow\infty$.
It will be demonstrated that for the matching mechanism and customer abandonment behavior considered in this study, this assumption holds true. Explicit formulae will be provided to support this assumption. The assumption that as $x_{ii}\rightarrow b_{i}$, $w_{i}^{d}(x_{ii})\rightarrow\infty$ arises from the customer behavior model, where customers exhibit impatience and will abandon the service if not served promptly. As a result, an infinite number of idle drivers is necessary to meet all customer demands. This aspect will be elaborated on further in subsequent sections. In what follows, for the sake of uniformity, we write the sojourn time as a function of $x$, i.e., $T_{ik}(x)$ for any $(i,k)\in S\times A$, even though it is a function of $x_{ik}$ if $i=k$ and a constant if $i\not=k$. Given Assumption 1, the mass rate satisfies the following equation for a stationary system,
(2) $\mu_{ik}=x_{ik}T_{ik}(x),\>\forall(i,k)\in S\times A.$
As the total mass is $m$, immediately, we obtain the conversation law of mass,
(3) $\sum_{(i,k)\in S\times A}x_{ik}T_{ik}(x)=\sum_{i\in S}x_{ii}\big{(}w_{i}^{d}(x_{ii})+\tau_{i}(x_{ii})+t_{i}\big{)}+\sum_{i,j\in S\atop i\not=j}x_{ij}t_{ij}=m.$
Driver’s Strategy. In this section, our focus is on the repositioning strategy of drivers. After completing a trip, drivers face a decision problem where they must choose between serving their current location or repositioning to another location that potentially offers more customers and higher income. We assume that every driver follows a stationary repositioning strategy, meaning that their choice of action is solely dependent on states (the current location) and is independent of time. A driver’s stationary strategy is characterized by the matrix $(\pi_{ik})_{i\in S,k\in A}$. Here, $\pi_{ik}$ represents the probability that a driver selects action $k$ where $\sum_{k\in A}\pi_{ik}=1$. Given a stationary mass rate $x$, the aggregate stationary repositioning strategy, which amalgamates the strategies of all drivers, can be derived as follows,
(4) $\pi_{ik}(x)=\frac{x_{ik}}{\sum_{k\in A}x_{ik}}$
for state $i$ with a positive denominator, while $\pi_{ik}=0$ otherwise. In fact, the mass rate encapsulates the full information regarding drivers’ strategies. As we will demonstrate later in Proposition 2.2, the optimality of drivers’ strategies is derived in terms of the vector of mass rates.
Reward Structure. In this model, we assume a normalization of zero operating costs for vehicles and zero opportunity costs for drivers. We do not address the participation problem for drivers and assume that all drivers are willing to join the platform and adhere to the matching arrangements set by the platform. We assume that rewards are continuously generated at a compensatory rate of $c_{i}>0$ solely during the time when a driver is transporting a customer from the customer’s pick-up location to their destination. Consequently, with each trip originating from region $i$, a driver can, on average, earn $c_{i}t_{i}$. Given a stationary system, the average reward per driver per unit of time, denoted by $\Phi$, is a function of stationary mass rate vector $x$ given by
(5) $\Phi(x)=\frac{1}{m}\sum_{i\in S}x_{ii}c_{i}t_{i}.$
This study focuses on the semi-Markov decision process with an average reward criterion where each driver aims to maximize their individual average reward over an infinite time horizon. Let $(\tilde{\pi}_{ik})_{i\in S,k\in A}$ be a reposition strategy for a representative driver. Let $\tilde{x}=(\tilde{x}_{ik})_{i\in S,k\in A}\in\mathbb{R}_{+}^{N^{2}}$ be the representative driver’s mass rate where for any $i$ and $k$, $\tilde{x}_{ik}$ represents the average number that this driver arrives in region $i$ and take action $k$ within one unit of time. Then $\tilde{\pi}$ and $\tilde{x}$ are related through a equation similar to (4),
(6) $\tilde{\pi}_{ik}=\frac{\tilde{x}_{ik}}{\sum_{k\in A}\tilde{x}_{ik}}$
Given the system’s aggregated stationary mass rate $x$, the representative driver’s $\tilde{x}$ satisfies the conversation law and balance equation as follows,
(7a) $\sum_{i\in S}\tilde{x}_{ii}\big{(}w_{i}^{d}(x_{ii})+\tau_{i}(x_{ii})+t_{i}\big{)}+\sum_{i,j\in S\atop i\not=j}\tilde{x}_{ij}t_{ij}=1.$
(7b) $\sum_{j\in S}\tilde{x}_{ij}=\sum_{{j\in S},j\not=i}\tilde{x}_{ji}+\sum_{j\in S}\tilde{x}_{jj}q_{ji},\>\forall i\in S$
The representative driver’s objective is to choose a reposition strategy $\tilde{\pi}$ to maximize his own average reward $\sum_{i\in S}\tilde{x}_{ii}c_{i}t_{i}$.
In summary, we have defined a non-atomic game, denoted by $\mathcal{G}$, through the tuple
$\left\langle m,S,A,\big{(}P^{k}_{ij}\big{)}_{i,j\in S,k\in A},\Big{(}T_{ij}(\cdot)\Big{)}_{i,j\in S},(c_{i})_{i\in A}\right\rangle.$
The game $\mathcal{G}$ belongs to a mean field type. A mean field game (MFG) is a game-theoretical framework that models the strategic decision-making of a large number of interacting agents (Jovanovic and Rosenthal, 1988; Huang et al., 2006; Hopenhayn, 1992; Adlakha et al., 2015). In MFGs, each agent interacts with the ‘mean field’ (in our context, the vector $\mu$ of agent mass per state-action pair), which represents the average effect of the strategies taken by all the other agents. In our model, a key mean field term is the sojourn time $\big{(}T_{ij}(\cdot)\big{)}_{i,j\in S}$, which implicitly depends on the ‘mean field’ $\mu$ through the mass rate vector $x$. In this paper we make no assumptions that the reader is familiar with MFG theory, and our analysis is self-contained.
Table 1. Notations and Units
  Notation Physical Meaning Suggested Unit
 $m$ Overall mass density of the drivers (vehicles) $\text{veh}/\text{km}^{2}$
$\mu_{ik}$ Mass density of the drivers in region $i$ taking action $k$ $\text{veh}/\text{km}^{2}$
$t_{ij}$ Expected travel time from region $i$ to $j$ min
$x_{ik}$ Rate of the drivers arrive in region $i$ and take action $k$ per unit area $\text{veh}/(\text{min}\cdot\text{km}^{2})$
$c_{i}$ Money paid to the driver per unit time when carrying a passenger \$$\$/\text{min}$
$b_{i}$ Mass of customers (passengers) arrive in region $i$ per unit time per unit area $\text{pass}/(\text{min}\cdot\text{km}^{2})$
$\theta_{i}$ Customer abandonment rate in region $i$ $\text{min}^{-1}$
$R_{i}$ Matching radius in region $i$ km
$w_{i}^{d}$ Driver’s expected waiting time in region $i$ min
$w_{i}^{c}$ Customer’s expected waiting time in region $i$ min
$\tau_{i}$ Expected pick-up time in region $i$ min
[…]
$\mu_{i}^{d}$ Mass density of the waiting drivers in region $i$ $\text{veh}/\text{km}^{2}$
$\mu_{i}^{c}$ Mass density of the waiting customers in region $i$ $\text{pass}/\text{km}^{2}$
Remark: We have not specified the size of each region in our setup. The region size will become relevant once we introduce the matching radius later in our analysis. Our methodology and analysis do not depend on the specific sizes of each individual region. It is more sensible to regard the mass we defined as mass per unit area (or mass density) and the rate as rate per unit area. Occasionally, we may omit the term “per unit area” when its context is apparent. To elucidate, we provide the physical meanings and suggested units for the primary variables and parameters in this paper, including those later introduced, in Table 1. These suggested units will be used in the numerical experiments and simulations conducted in this paper.
2.2. Stationary Equilibrium
Now we can give the definition of a stationary equilibrium of the game $\mathcal{G}$.
Definition 2.1.
A stationary equilibrium of game $\mathcal{G}$ is a vector $(\mu^{\dagger},x^{\dagger})\in\mathcal{U}_{m}(S\times A)\times\mathbb{R}_{+}^{N^{2}}$ such that
  • (a)
    $x^{\dagger}$ is the mass rate under the mass distribution $\mu^{\dagger}$ and it is stationary, i.e., $(\mu^{\dagger},x^{\dagger})$ satisfies (1), (2), and (3).
  • (b)
    Given the aggregate behavior of all the drivers characterized by $(\mu^{\dagger},x^{\dagger})$, the average reward per unit mass $\Phi(x^{\dagger})$ in (5) is the optimal average reward for a representative driver. In other words, the behavioral strategy (aggregated strategy) $\pi(x^{\dagger})$ in (4) is the best response against $(\mu^{\dagger},x^{\dagger})$.
The second condition states that no driver can achieve a higher average reward by deviating from the current aggregated behavioral strategy. A stationary equilibrium of the game can be fully characterized by the mass rate vector, as stated in the following proposition.
Proposition 2.2.
$x^{\dagger}$ is an equilibrium mass rate of game $\mathcal{G}$ if and only if $x^{\dagger}$ is an optimal solution of the following optimization problem,
(8) $\max\quad$ $\sum_{i\in S}x_{ii}c_{i}t_{i}$
(8a) s.t. $\sum_{i\in S}x_{ii}\big{(}w_{i}(x^{\dagger}_{ii})+\tau_{i}(x^{\dagger}_{ii})+t_{i}\big{)}+\sum_{i,j\in S\atop i\not=j}x_{ij}t_{ij}=m,$
(8b) $\sum_{j\in S}x_{ij}=\sum_{j\in S,j\not=i}x_{ji}+\sum_{j\in S}q_{ji}x_{jj},\>i\in S,$
(8c) over $x=(x_{ij})_{i,j\in S}\in\mathbb{R}_{+}^{N^{2}}.$
Proof.
Consider a mass rate $x^{\dagger}\in\mathbb{R}_{+}^{N^{2}}$, we will identify the conditions for $x^{\dagger}$ being an equilibrium mass rate. Note that we can compute the corresponding distribution $\mu^{\dagger}$ by (2). Hence, the first condition for $x^{\dagger}$ being an equilibrium mass rate, namely, Definition 2.1(a), is equivalent to that $x^{\dagger}$ satisfies (8a) and (8b) as per (3) and (1) respectively. Let $\tilde{x}=(\tilde{x}_{ik})_{i\in S,k\in A}\in\mathbb{R}_{+}^{N^{2}}$ be the representative driver’s mass rate. Since the representative driver’s strategy is encoded in $\tilde{x}$ through (6), he seeks to manipulate $\tilde{x}$ to maximize his average reward, i.e. $\sum_{i\in S}\tilde{x}_{ii}c_{i}t_{i}$. A feasible $\tilde{x}$ must satisfies (7b) in a stationary system. Moreover, as the sojourn time is determined by the aggregate mass rate $x^{\dagger}$, similar to (7a), a feasible $\tilde{x}$ must satisfy that
$\sum_{i\in S}\tilde{x}_{ii}\big{(}w_{i}(x^{\dagger}_{ii})+\tau_{i}(x^{\dagger}_{ii})+t_{i}\big{)}+\sum_{i,j\in S\atop i\not=j}\tilde{x}_{ij}t_{ij}=1.$
Thus, the representative driver’s optimization problem is,
(9) $\max\quad$ $\sum_{i\in S}\tilde{x}_{ii}c_{i}t_{i}$
(9a) s.t. $\sum_{i\in S}\tilde{x}_{ii}\big{(}w_{i}(x^{\dagger}_{ii})+\tau_{i}(x^{\dagger}_{ii})+t_{i}\big{)}+\sum_{i,j\in S\atop i\not=j}\tilde{x}_{ij}t_{ij}=1,$
(9b) $\sum_{j\in S}\tilde{x}_{ij}=\sum_{j\in S,j\not=i}\tilde{x}_{ji}+\sum_{j\in S}q_{ji}\tilde{x}_{jj},\>i\in S,$
(9c) over $\tilde{x}=(\tilde{x}_{ij})_{i,j\in S}\in\mathbb{R}_{+}^{N^{2}}.$
The second condition for $x^{\dagger}$ being an equilibrium mass rate, namely, Definition 2.1(b), is equivalent to that $\tilde{x}=\frac{1}{m}x^{\dagger}$ solves the optimization problem (9). Therefore, the two conditions for $x^{\dagger}$ being an equilibrium mass rate, i.e., Definition 2.1(a) and Definition 2.1(b) together, are equivalent to that $x^{\dagger}$ solves the optimization problem (8). ∎
Proposition 2.2 provides a way to verify whether a given $x^{\dagger}$ is an equilibrium mass rate of game $\mathcal{G}$. We first compute the sojourn times $w_{i}(x^{\dagger}_{ii})$ and $\tau_{i}(x^{\dagger}_{ii})$ to obtain the linear constraints (8a) and then check if $x^{\dagger}$ solves the linear optimization problem (8).
A key result of this paper is the establishment of a one-to-one correspondence between the stationary equilibria and the Karush-Kuhn-Tucker (KKT) points of an optimization problem. Note that for a general optimization problem, KKT points refer to those points that satisfy the KKT conditions.
Theorem 2.3.
Suppose Assumption 1 is true, an $x^{\dagger}\in\mathbb{R}_{+}^{N^{2}}$ is an equilibrium mass rate if and only if $x^{\dagger}$ satisfies the KKT conditions of the optimization problem:
(10) $\max\quad$ $m\log\bigg{(}\sum_{i\in S}x_{ii}c_{i}t_{i}\bigg{)}-\sum_{i\in S}x_{ii}t_{i}-\sum_{i,j\in\mathcal{S}\atop i\not=j}x_{ij}t_{ij}-\sum_{i\in S}\int_{0}^{x_{ii}}w_{i}^{d}(s)\mathrm{d}s-\sum_{i\in S}\int_{0}^{x_{ii}}\tau_{i}(s)\mathrm{d}s$
s.t. $\sum_{j\in S}x_{ij}=\sum_{j\in S,j\not=i}x_{ji}+\sum_{j\in S}q_{ji}x_{jj},\>i\in\mathcal{S},$
over $x=(x_{ij})_{i,j\in S}\in\mathbb{R}_{+}^{N^{2}}.$
Proof.
Let $x^{\dagger}$ be an equilibrium. By Proposition 2.2, $x^{\dagger}$ is an optimal solution of (8). $x^{\dagger}$ remains optimal if the objective in (8) is replaced by
(11) $m\log\bigg{(}\sum_{i}x_{ii}c_{i}t_{i}\bigg{)}.$
Thus, $x^{\dagger}$ satisfies the first order conditions:
(12a) $\frac{mc_{i}t_{i}}{\sum_{i}x^{\dagger}_{ii}c_{i}t_{i}}-\lambda\bigg{[}t_{i}+\tau_{i}(x^{\dagger}_{ii})+w_{i}^{d}(x^{\dagger}_{ii})\bigg{]}+\sum_{j\not=i}(\nu_{i}-\nu_{j})q_{ij}\leq 0,\text{ with equality if }x^{\dagger}_{ii}>0,\ \forall i,$
(12b) $-\lambda t_{ij}+\nu_{i}-\nu_{j}\leq 0,\text{ with equality if }x^{\dagger}_{ij}>0,\ \forall i,j,i\not=j,$
where $\lambda\in\mathbb{R}$ is the Lagrange multiplier of the conservation law (8a) and $\nu_{i}\in\mathbb{R}$ is the Lagrange multiplier of the $i$th balance equation (8b).
By multiplying both sides of (12a) and (12b) by the respective $x^{\dagger}_{ij}$ for all $i$, $j$, then summing them up, we obtain:
$\sum_{i}x^{\dagger}_{ii}\frac{mc_{i}t_{i}}{\sum_{i}x^{\dagger}_{ii}c_{i}t_{i}}-\lambda\sum_{i}x^{\dagger}_{ii}\bigg{[}t_{i}+\tau_{i}(x^{\dagger}_{ii})+w_{i}^{d}(x^{\dagger}_{ii})\bigg{]}-\lambda\sum_{j\not=i}x^{\dagger}_{ij}t_{ij}$
$+\sum_{i}x^{\dagger}_{ii}\sum_{j\not=i}(\nu_{i}-\nu_{j})q_{ij}+\sum_{j\not=i}(\nu_{i}-\nu_{j})x^{\dagger}_{ij}=0$
Apply the conservation law (8a) and balance equation (8b), we obtain $m(1-\lambda)=0$, so $\lambda=1$. Therefore, the first order conditions can be rewritten as
(13a) $\frac{mc_{i}t_{i}}{\sum_{i}x^{\dagger}_{ii}c_{i}t_{i}}-t_{i}-\tau_{i}(x^{\dagger}_{ii})-w_{i}^{d}(x^{\dagger}_{ii})+\sum_{j\not=i}(\nu_{i}-\nu_{j})q_{ij}\leq 0,\text{ with equality if }x^{\dagger}_{ii}>0,\ \forall i$
(13b) $-t_{ij}+\nu_{i}-\nu_{j}\leq 0,\text{ with equality if }x^{\dagger}_{ij}>0,\ \forall i,j,i\not=j$
where $\nu_{i}$ is the Lagrange multiplier of the $i$th balance equation. The conditions (13a) and (13b) are exactly the KKT conditions for the optimization problem (10).
Conversely, if $x^{\dagger}$ satisfies the KKT conditions of the optimization problem (10), i.e., (13a) and (13b), it also satisfies the first order conditions (12a) and (12b) for optimization problem (8) with $\lambda=1$ and objective function replaced by (11) . By multiplying both sides of (13a) and (13b) by the respective $x^{\dagger}_{ij}$ for all $i$, $j$, then summing them up, we obtain the conservation law (8a). Since $x^{\dagger}$ also satisfies the balance equation (8b), $x^{\dagger}$ is feasible and hence optimal for optimization problem (8). Thus, $x^{\dagger}$ is an equilibrium. ∎
Corollary 2.4.
If Assumption 1 holds, a stationary equilibrium always exists for game $\mathcal{G}$.
Proof.
It suffices to show that there exists at least one optimal solution to (10). This is because optimization problem (10) has linear constraints and hence, any optimal solution must adhere to the Karush–Kuhn–Tucker conditions. Then the existence of an equilibrium follows from Theorem 2.3. The current domain for variable $x$ in (10), $\mathbb{R}_{+}^{N^{2}}$, is open and unbounded. We need to show that the optimal value of (10) can be obtained in a closed and bounded (compact) subset of $\mathbb{R}_{+}^{N^{2}}$.
First, it should be noted that the optimal value of (10) must be obtained away from the zero vector, which is not a feasible solution. This is because as $||x||\rightarrow 0$, the objective value tends towards negative infinity. Second, as per Assumption 1, the mass rate $x_{ii}$ has an strict upper bound $b_{i}$ for any $i$. Third, for any $i\not=j$, as $x_{ij}$ goes to infinity, the objective function in (10) goes to negative infinity since $t_{ij}>0$. Consequently, the optimal value of (10) must be obtained when any $x_{ij}$ with $i\not=j$ is finite. Furthermore, the optimal value must be obtained when $x_{ii}$ is strictly less than $b_{i}$ for any $i$ since the derivative of the objective function in (10) with respect to $x_{ii}$,
$\frac{mc_{i}t_{i}}{\sum_{i}x^{\dagger}_{ii}c_{i}t_{i}}-t_{i}-\tau_{i}(x^{\dagger}_{ii})-w_{i}^{d}(x^{\dagger}_{ii})$
goes to negative infinity as $x_{ii}$ approaches $b_{i}$. Thus, without loss of generality, the open domain $\mathbb{R}_{+}^{N^{2}}$ for $x$ can be replaced by a closed and bounded region in $\mathbb{R}_{+}^{N^{2}}$. Given the continuity of the objective function, the optimal problem (10) must possess an optimal solution in $\mathbb{R}_{+}^{N^{2}}$. ∎
The optimization problem is convex when the (total) sojourn time $T_{ii}(x_{ii})$ increases with respect to $x_{ii}$ for every $i$. Moreover, if this increases strictly for every $i$, the optimization problem becomes strictly convex, leading to uniqueness of the stationary equilibrium.
Corollary 2.5.
Suppose i) Assumption 1 holds, and ii) the total sojourn time of the driver, $T_{i}(x_{ii})$, is a strictly increasing function of $x_{ii}$ for every $i\in S$. Then, a unique stationary equilibrium exists for the game $\mathcal{G}$.
On the other hand, if $T_{i}(x_{ii})$ is not an increasing function of $x_{ii}$, multiple equilibria can arise. In fact, under most dispatch algorithms, the pick-up time is not a monotonic function of $x_{ii}$. This non-monotonicity leads to the “Wild Goose Chase” equilibrium, where drivers may be dispatched to pick up customers located far away. This issue has also been investigated in prior work (Castillo et al., 2023). We will provide a more detailed mathematical analysis of WGC to investigate the impact of dispatch algorithms. We propose a novel dispatch algorithm designed to ensure that $T_{i}(x_{ii})$ is an increasing function of $x_{ii}$. This guarantees a unique equilibrium and eliminates WGC equilibrium.
3. Ride-Matching Algorithm under Abandonment Behavior
In this section, we present a comprehensive analysis of the platform’s ride-matching algorithms and customer abandonment behavior, examining their influence on the equilibrium and overall performance of the ride-hailing system. We show that the widely used nearest-neighbor dispatch, either without a matching radius or with an arbitrarily chosen one, can result in “Wild Goose Chase” equilibrium and the emergence of multiple equilibria. This underscores the importance of optimizing the design of the matching radius, a topic we explore in the next section.
3.1. Customer: Abandonment Behavior
Ride-hailing platforms typically permit customers to cancel the service if they consider their wait time to be excessively long. Our model, similar to (Yang et al., 2020; Wang et al., 2022), assumes that each customer has a maximum willing waiting time. If a customer is not matched with a driver within this time threshold, they will exit the system. Additionally, we assume that once a match is made, customers are committed to the ride and cannot abandon the service.
Suppose that each waiting customer in region $i$ has an exponential clock with parameter $\theta_{i}$ that if it expires prior to the customer being matched with a driver, this customer abandons the system. In the spirit of mean-field approximation, if at any time the mass density of waiting customer is $\mu_{i}^{c}$, these will generate a rate of abandonment $\theta_{i}\mu_{i}^{c}$, where
(14) $\theta_{i}\mu_{i}^{c}+x_{ii}=b_{i}$
i.e., abandonment rate + service rate (driver supply rate) = arrival rate. This equation establishes a relationship between the unknown variables $x_{ii}$, $\mu_{i}^{c}$ and known parameters $b_{i}$, $\theta_{i}$ using mean-field approximation.
3.2. Platform: Matching Policy
Figure 1. Two-matching-radius Nearest-neighbor Dispatch: (a). an arriving customer is immediately matched with the closest idle driver within matching radius $R_{c}$ and (b). an arrival driver is immediately matched with the closest waiting customer within matching radius $R_{d}$.
We assume the platform adopt a notable dispatch algorithm: nearest-neighbor dispatch (NN), which is simple and near-optimal (Besbes et al., 2022). Recent work propose using a matching radius to prevent drivers from being dispatched to pick up customers who are too far away, thereby improve the system performance (Castillo et al., 2023; Yang et al., 2020). Our observation is that only one matching radius is lack of flexibility. A key innovation of this study is the introduction of a two-matching-radius nearest-neighbor dispatch algorithm:
  • When a new customer arrives (as in Fig. 1.(a) ), the platform assigns him to the closest available driver within the first matching radius ($R_{c}$).
  • When a driver becomes available (as in Fig. 1.(b) ), the platform assigns him to the closest waiting customer within a second matching radius ($R_{d}$), which may differ from the first matching radius.
Drivers who are not immediately matched upon arrival will wait until they are paired with incoming customers. Similarly, customers who are not instantly matched upon arrival will either wait to be paired with incoming drivers or abandon the service if they become impatient.
3.3. Formulas of Waiting Time and Pick-up Time
In this section we establish formulas for the average waiting time and pick-up time as claimed in Assumption 1 based on the customer abandonment behavior defined in Section 3.1 and the platform’s dispatching algorithm defined in Section 3.2.
Let each region $i$’s area be denoted by $A_{i}$. We introduce two matching radii in region $i$, $R_{i,c},R_{i,d}\in(0,\sqrt{A_{i}/\pi})$, where $\sqrt{A_{i}/\pi}$ is the approximate radius of region $i$. We denote the average vehicle speed by $v_{i}$. Let $\mu_{i}^{c}$ and $\mu_{i}^{d}$ denote the average densities of waiting customers and waiting drivers respectively across the region $i$. Recall that $x_{ii}$ denotes rate of drivers that become available and choose to serve region $i$ per unit area, and $b_{i}$, $\theta_{i}$ the arrival rate, abandonment rate of customers per unit area, respectively.
We assume the locations of waiting customers and waiting drivers are drawn from independent uniform distributions across region $i$ at any moment. This assumption is justified by the randomness of customers’ origins and destinations, as well as the mobility of the vehicles. Given the matching radius $R_{i,c}$, when a driver becomes available in region $i$ and decides to provide service in this region, the probability that the driver is within distance $R_{i,c}$ of a randomly chosen waiting customer in region $i$ is $\tfrac{\pi R_{i,c}^{2}}{A_{i}}$. Assume there are $N_{i}^{c}$ waiting customers in region $i$ at this moment, then the driver is not matched with any of these waiting customers with probability
$(1-\frac{\pi R_{i,c}^{2}}{A_{i}})^{N_{i}^{c}}.$
Similarly, given the matching radius $R_{i,d}$, if there are $N_{i}^{d}$ waiting drivers in region $i$, a new arrival customer is not matched with any of these waiting customers with probability
$(1-\frac{\pi R_{i,d}^{2}}{A_{i}})^{N_{i}^{d}}.$
Now we can find the formula for driver’s average waiting time. A key observation is that for drivers, the rate at which they enter the waiting status equals the rate at which they exit it. This exit rate is equivalent to the rate at which arriving customers are matched with drivers. We apply mean-field approximation to replace stochastic variables, such as $N_{i}^{c}$ and $N_{i}^{d}$, with their expected values: $\mu_{i}^{c}A_{i}$ and $\mu_{i}^{d}A_{i}$ respectively. Then, in steady state, the relation between $\mu_{i}^{c}$ and $\mu_{i}^{d}$ is
(15) $x_{ii}(1-\frac{\pi R_{i,c}^{2}}{A_{i}})^{\mu_{i}^{c}A_{i}}+b_{i}(1-\frac{\pi R_{i,d}^{2}}{A_{i}})^{\mu_{i}^{d}A_{i}}=b_{i}$
Note that $\mu_{i}^{c}$ and $x_{ii}$ satisfy (14), $\mu_{i}^{c}$ can be expressed as a function of $x_{ii}$,
(16) $\mu_{i}^{c}(x_{ii})=\frac{1}{\theta}(b_{i}-x_{ii}).$
As a result, by (15), the unknown variable $\mu_{i}^{d}$ can also be expressed a as a function of $x_{ii}$,
(17) $\mu_{i}^{d}(x_{ii})=\frac{\log(1-\frac{x_{ii}}{b_{i}}(1-\frac{\pi R_{i,c}^{2}}{A_{i}})^{\frac{b_{i}-x_{ii}}{\theta_{i}}A_{i}})}{A_{i}\log(1-\frac{\pi R_{i,d}^{2}}{A_{i}})}\,.$
Let $w_{i}^{d}$ denote the average waiting time of the drivers. By Little’s law, $\mu_{i}^{d}=x_{ii}w_{i}^{d}$. Hence, $w_{i}^{d}$ can be expressed as a function of $x_{ii}$,
(18) $w_{i}^{d}(x_{ii})=\frac{\log(1-\frac{x_{ii}}{b_{i}}(1-\frac{\pi R_{i,c}^{2}}{A_{i}})^{\frac{b_{i}-x_{ii}}{\theta_{i}}A_{i}})}{x_{ii}A_{i}\log(1-\frac{\pi R_{i,d}^{2}}{A_{i}})}\,.$
One can check that $w_{i}^{d}(x_{ii})\rightarrow\infty$ as $x_{ii}\rightarrow b_{i}$ as we claimed in Assumption 1.
Next, we derive the formula for expected pick-up time, which is amount to finding the expected pick-up distance. We compute this in two distinct cases:
  • Matched with no waiting. In this case, a driver arrives and is successfully matched with a waiting customer. The probability of this occurring is
    $1-(1-\frac{\pi R_{i,c}^{2}}{A_{i}})^{\mu_{i}^{c}A_{i}}.$
    If matched, the probability that the pick-up distance is less than $r$ is the probability that at least one waiting customer is within distance $r\in(0,R_{i,c}]$ of the driver, which is
    $1-(1-\frac{\pi r^{2}}{A_{i}})^{\mu_{i}^{c}A_{i}}$
    Consequently, the conditional expectation of the pick-up distance is
    $\frac{1}{1-(1-\frac{\pi R_{i,c}^{2}}{A_{i}})^{\mu_{i}^{c}A_{i}}}\int_{0}^{R_{i,c}}r\mathrm{d}\Big{(}1-(1-\frac{\pi r^{2}}{A_{i}})^{\mu_{i}^{c}A_{i}}\Big{)}$
  • Matched with waiting. In this case, a driver arrives and must wait until they are matched with a newly arriving customer. The probability of this occurring is,
    $(1-\frac{\pi R_{i,c}^{2}}{A_{i}})^{\mu_{i}^{c}A_{i}}.$
    The expected pick-up distance can be computed from the perspective of the arriving customer. The probability that a customer is successfully matched with a waiting driver is,
    $1-(1-\frac{\pi R_{i,d}^{2}}{A_{i}})^{\mu_{i}^{d}A_{i}}.$
    The probability that the pick-up distance is less than $r$ is the probability that at least one waiting driver is within distance $r\in(0,R_{i,d}]$ of the customer, which is
    $1-(1-\frac{\pi r^{2}}{A_{i}})^{\mu_{i}^{d}A_{i}}$
    Consequently, the conditional expectation of the pick-up distance is
    $\frac{1}{1-(1-\frac{\pi R_{i,d}^{2}}{A_{i}})^{\mu_{i}^{d}A_{i}}}\int_{0}^{R_{i,d}}r\mathrm{d}\Big{(}1-(1-\frac{\pi r^{2}}{A_{i}})^{\mu_{i}^{d}A_{i}}\Big{)}$
In summary, the expected pick-up time is a function of $x_{ii}$ as follows,
(19) $\tau_{i}(x_{ii})=\frac{1}{v_{i}}\int_{0}^{R_{i,c}}(1-\frac{\pi r^{2}}{A_{i}})^{\mu_{i}^{c}A_{i}}-(1-\frac{\pi R_{i,c}^{2}}{A_{i}})^{\mu_{i}^{c}A_{i}}\mathrm{d}r+\frac{b_{i}}{v_{i}x_{ii}}\int_{0}^{R_{i,d}}(1-\frac{\pi r^{2}}{A_{i}})^{\mu_{i}^{d}A_{i}}-(1-\frac{\pi R_{i,d}^{2}}{A_{i}})^{\mu_{i}^{d}A_{i}}\mathrm{d}r,$
where $\mu_{i}^{c}$ and $\mu_{i}^{d}$ are functions of $x_{ii}$ as in (16) and (17) respectively.
Approximate Formulas. As the formulas for waiting time and pick-up time are too complex, we derive approximate formulas below. We apply a change of variable trick,
(20) $-\frac{A_{i}}{\pi}\log(1-\frac{\pi R_{i,\cdot}^{2}}{A_{i}})\mapsto R_{i,\cdot}^{2},$
where we keep the same notation for the new radius with a slight abuse of notation. Then, (15) can be rewritten as,
$x_{ii}e^{-\mu_{i}^{c}\pi R_{i,c}^{2}}+b_{i}e^{-\mu_{i}^{d}\pi R_{i,d}^{2}}=b_{i}.$
Furthermore, the drivers’ waiting times and pick-up times can be approximated as:
(21) $\begin{cases}w_{i}^{d}(x_{ii})&=\frac{-\log(1-\frac{x_{ii}}{b_{i}}e^{-\frac{b_{i}-x_{ii}}{\theta_{i}}\pi R_{i,c}^{2}})}{x_{ii}\pi R_{i,d}^{2}}\\ \tau_{i}(x_{ii})&=\frac{1}{v_{i}}\int_{0}^{R_{i,c}}e^{-\mu_{i}^{c}(x_{ii})\pi r^{2}}-e^{-\mu_{i}^{c}(x_{ii})\pi R_{i,c}^{2}}\mathrm{d}r+\frac{b_{i}}{v_{i}x_{ii}}\int_{0}^{R_{i,d}}e^{-\mu_{i}^{d}(x_{ii})\pi r^{2}}-e^{-\mu_{i}^{d}(x_{ii})\pi R_{i,d}^{2}}\mathrm{d}r\end{cases}$
Note that the left hand side of (20) is approximately $R_{i,\cdot}^{2}$ when $R_{i,\cdot}<<\sqrt{A_{i}}$ as $\log(1-y)\approx-y$ when $y$ approaches 0 and goes to $\infty$ as $\pi R_{i,\cdot}^{2}\rightarrow A_{i}$. Hence, when calculating the pick-up time we do not multiply the change of variable factor when computing the integral. Numerical results show that this is still a good estimation of the original pick-up time even when the original radius $R_{i}$ is large. There is an alternative point of view which yields the same approximate formula in (21): assume each region $i$ is an unbounded plane with matching radii $R_{i,c},R_{i,d}\in(0,\infty)$. In this unbounded plane, the locations of waiting customers and drivers follow spatial Poisson processes, where $\mu_{i}^{c}$ and $\mu_{i}^{d}$ are the waiting customer and waiting driver intensity parameters, respectively. In fact, the change of variable in equation (20) defines a map from the original finite region to an infinite region. Although this transformation distorts the Euclidean metric in the original region, the distortion is negligible when the matching radius is sufficiently small. For large matching radii, while the distortion cannot be entirely ignored, the introduced approximation error remains acceptable according to numerical results.
3.4. Analysis of Equilibria in One Region: Wild Goose Chase
In this section, we show that the original NN dispatch algorithm, where the matching radius $R_{i,\cdot}$ is set to $\sqrt{A_{i}/\pi}$ (or $\infty$ if the approximated formulas (21) are used), can still lead to multiple equilibria even when there is only one region. To simplify notation, we omit the subscript $i$ since there is only one region. As there is no driver repositioning, $x$ denotes the driver supply rate. When the two matching radii are $\sqrt{A/\pi}$, the equation (15) becomes,
$x\cdot 0^{\mu^{c}}+b\cdot 0^{\mu^{d}}=b$
which seems to be meaningless. But if we define the indeterminate limit [^2] Note that $\lim\limits_{x\rightarrow 0}\lim\limits_{y\rightarrow 0}x^{y}=1$ and $\lim\limits_{y\rightarrow 0}\lim\limits_{x\rightarrow 0}x^{y}=0$. $0^{0}$ to be $1$, we obtain two possible solutions,
$\mu^{c}=0,\mu^{d}>0\text{ or }\mu^{c}>0,\mu^{d}=0.$
These solutions are validated through an intuitive argument: it is impossible for both the number of waiting customers and the number of waiting drivers to be positive when the matching radius is infinite, as they would be matched with each other. Consequently, there are two distinct types of equilibriua:
  • Oversupply Equilibrium: There is an oversupply of drivers, who must wait to be matched with customers. This results in a positive average waiting time for drivers.
  • Undersupply Equilibrium: There is an undersupply of drivers, leading to customers waiting for available drivers. This results in a positive average waiting time for customers.
Before presenting the equilibrium results, we must first discuss how to calculate the pick-up time using Equation (19) or the approximate formula in Equation (21) under this extreme case.
Approximate Pick-up Time. When $R_{c}=R_{d}=\sqrt{A/\pi}$ (or $\infty$), our formulae for $\tau_{i}$ in (19) and (21) can be still used to estimate the pick-up time. In fact, both formulas give the same asymptotical result. We assume $\mu^{c}=0,\mu^{d}>0$. In this case, $x=b$ as $\mu^{c}=0$. Applying (19), the average pick-up time is estimated by
$\tau(x)=\frac{1}{v}\int_{0}^{\sqrt{\frac{A}{\pi}}}(1-\frac{\pi r^{2}}{A})^{\mu^{d}A}\mathrm{d}r\overset{\frac{\pi r^{2}}{A}\mapsto s^{2}}{=}\frac{1}{v}\sqrt{\frac{A}{\pi}}\int_{0}^{1}(1-s^{2})^{\mu^{d}A}\mathrm{d}s=\left(\frac{1}{\sqrt{\pi}}\int_{0}^{1}(1-s^{2})^{\mu^{d}A}\mathrm{d}s\right)\frac{\sqrt{A}}{v}.$
Using the change of variable trick (i.e., (21)), the average pick-up time is
(22) $\tau(x)=\frac{1}{v}\int_{0}^{\infty}e^{-\mu^{d}\pi r^{2}}\mathrm{d}r=\frac{1}{2v\sqrt{\mu^{d}}}=\frac{1}{2\sqrt{\mu^{d}A}}\frac{\sqrt{A}}{v},$
which is exactly the famous square root law in literature (e.g., see (Besbes et al., 2022)). Recall that $\sqrt{\frac{A}{\pi}}$ estimates the radius of the whole region and $\mu^{d}A$ is the mean of the number of waiting drivers ($N^{d}$) in the region. We can numerically check that for $n\in\mathbb{N}^{+}$,
$\left(\frac{1}{\sqrt{\pi}}\int_{0}^{1}(1-s^{2})^{n}\mathrm{d}s\right)-\frac{1}{2\sqrt{n}}$
is small even for small $n$ and goes to 0 as $n\rightarrow\infty$. In fact, for $y\in(0,\infty)$
$\int_{0}^{1}(1-s^{2})^{y}\mathrm{d}s\overset{s^{2}\mapsto s}{=}\frac{1}{2}\int_{0}^{1}s^{-\frac{1}{2}}(1-s)^{y}\mathrm{d}s=\frac{1}{2}B(\frac{1}{2},y+1)=\frac{1}{2}\frac{\Gamma(\frac{1}{2})\Gamma(y+1)}{\Gamma(y+\frac{3}{2})}\overset{y\rightarrow\infty}{\rightarrow}\frac{\sqrt{\pi}}{2\sqrt{y}},$
where $B(y_{1},y_{2})=\int_{0}^{1}s^{y_{1}-\frac{1}{2}}(1-s)^{y_{2}-1}\mathrm{d}t$ is the Euler’s beta function and $\Gamma(y)=\int_{0}^{\infty}s^{y-1}e^{-s}\mathrm{d}s$ is the gamma function. Similarly, when $\mu^{c}>0,\mu^{d}=0$, $\tau(x)\rightarrow\tfrac{1}{2v\sqrt{\mu^{c}}}$ if $\mu^{c}A=\frac{b-x}{\theta}A\rightarrow\infty$. Therefore, we will use the square root law (22) to estimate the pick-up time.
We analyze the two types of equilibria as follows.
Case 1: Oversupply Equilibrium. In this case, there is an oversupply of drivers, and hence a positive density of waiting drivers, i.e., $\mu^{c}=0,\mu^{d}>0$. Given the pick-up time (22), the conversation law of mass (3) gives,
$\mu^{d}+\frac{b}{2v\sqrt{\mu^{d}}}+bt=m$
which has one solution of equilibrium when $m=\tfrac{1}{3}\left(\tfrac{b^{2}}{16v^{2}}\right)^{\frac{1}{3}}+bt$, and two solutions of equilibria when $m>\tfrac{1}{3}\left(\tfrac{b^{2}}{16v^{2}}\right)^{\frac{1}{3}}+bt.$ This type of equilibria have the same driver supply rate. We denote the equilibrium driver supply rate in this case by $x^{o}$ and then $x^{o}=b$ since every customer is served in this oversupply regime.
Case 2: Undersupply Equilibrium. In this case, there is an oversupply of drivers, and hence a positive density of waiting customers, i.e., $\mu^{c}>0,\mu^{d}=0$. Then $w^{d}=0$ as $\mu^{d}=0$. The pick-up time is given by a square root law similar to (22), and the conversation law of mass (3) becomes,
(23) $\frac{x}{2v\sqrt{\mu^{c}}}+xt=\frac{x}{2v}\sqrt{\frac{\theta}{b-x}}+xt=m$
which has at least one solution of equilibrium as long as $m>0$. We denote the equilibrium service rate in this case by $x^{u}$ where $x^{u}$ solves equation (23). Here, $x^{u}<b$, which is a “bad” equilibrium compared to the oversupply regime. The equation (23) may admit solutions with extremely small driver supply rates and large pick-up times.
Infinity Price of Anarchy. The undersupply equilibrium can be arbitrarily bad compared to the oversupply equilibrium, i.e., $x^{u}<<b$. Specifically, when the mass is $m=\tfrac{1}{3}\left(\tfrac{b^{2}}{16v^{2}}\right)^{\frac{1}{3}}+bt$, there are exactly two equilibria The two equilibria are $x^{o}=b$ and $x^{u}$ where $\frac{x^{u}}{b}$ satisfies
$\frac{x^{u}}{b}=\frac{\frac{1}{3}\Big{(}\frac{b^{2}}{16v^{2}}\Big{)}^{\frac{1}{3}}+bt}{\frac{b}{2v}\sqrt{\frac{\theta}{b-x^{u}}}+bt},$
which approaches 00 if $\theta b^{3}=1$ and $b\rightarrow 0$. Therefore, the “bad” equilibrium ($x^{u}$) can be arbitrarily worse than the “good” equilibrium ($x^{o}$). Furthermore, an infinity price of anarchy occurs under the following conditions: 1) the region size is sufficiently large ($A\rightarrow\infty$); 2) the customer arrival rate is extremely low ($b\rightarrow 0$); and 3) the customers exhibit extreme impatience ($\theta\rightarrow\infty$). This extreme scenario is primarily theoretical and rarely observed in practice. However, it highlights a critical insight: if the dispatch algorithm is not properly designed, the system’s efficiency loss can be catastrophic.
Wild Goose Chase. The WGC equilibrium exists in the undersupply regime (characterized by the “bad” equilibrium $x^{u}$), where drivers are assigned to pick up customers located far away. In this regime, the system experiences a driver shortage because: 1) high average pick-up times force drivers to travel long distances to reach customers; 2) wasted time en route further reduces driver availability, exacerbating the undersupply. This creates a vicious cycle: longer travel times lead to fewer drivers being available to respond to new requests, worsening the imbalance between supply and demand.
Table 2. Values of Parameters
 Parameters $A$ $b$ $v$ $\frac{1}{\theta}$ $t$ $m$ $R$
[…]
 Values $100$ $2$ $0.5$ $1$ $10$ $(0,\infty)$$\{1,2,3,4\}$
Figure 2. The driver density (RHS of (24)) as a function of the driver supply rate $x$. We set the matching radius $R_{c}$ and $R_{d}$ to be same. In this figure, from left to right, the value of matching radius is 0.5 km, 1 km, 2 km, and 3 km.
Impact of Matching Radius. The use of fixed matching radius with $R_{c},R_{d}<\sqrt{A/\pi}$, does not completely eliminate multiple equilibria or prevent the WGC equilibrium. In the one region case. we can compute the equilibria by employing the conversation law of mass in (3), i.e.,
(24) $x\big{(}w^{d}(x)+\tau(x)+t\big{)}=m,$
where $w^{d}(x)$ and $\tau(x)$ are given in (18) and (19) . Since $\tau(x)$ is non-monotonic, multiple equilibria can arise for specific parameter choices.
[…]
Extension to Multiple Region. When multiple regions are considered, the average pick-up time in each region is a non-monotonic function of the local driver supply rate, leading to the potential for even more equilibria. In the single-region case, there can be three equilibria. As the number of regions increases, the number of equilibria grows combinatorially. We omit the detailed analysis here due to its complexity, but this can be verified using a toy model, such as a spoke-and-hub network.
4. Dynamically Optimize Matching Radii
In this section, we demonstrate how to address the “Wild Goose Chase” problem and eliminate multiple equilibria by implementing an optimal dynamic adjustment of matching radii.
The multiple equilibria and WGC are not resolved by introducing fixed matching radii. Instead of treating matching radii as fixed parameters of the dispatch algorithm, we propose to dynamically optimize matching radii. Specifically, we treat the waiting time $w_{i}^{d}$ and pick-up time $\tau_{i}$ in (21) as functions of both the driver supply rate $x_{ii}$ and the matching radii $R_{i,c}$ and $R_{i,d}$. Our goal is to find the optimal radii $R_{i,c}$ and $R_{i,d}$ that minimize the sum of waiting and pick-up times:
(25) $(R_{i,c}(x_{ii}),R_{i,c}(x_{ii}))=\arg\min_{(R_{i,c},R_{i,d})\in\mathbb{R}^{2}_{+}}w_{i}^{d}(x_{ii},R_{i,c},R_{i,d})+\tau_{i}(x_{ii},R_{i,c},R_{i,d}),\quad\forall i\in S,$
where $w_{i}^{d}$ and $\tau_{i}$ are given by approximation formulas in (21) since the original formulas (18) and (19) we derive can be approximated by (21) when $R_{c},R_{d}$ are small. We will show that the above optimization problem in (25) can be solved and the dynamic matching radii function $(R_{i,c}(x_{ii}),R_{i,d}(x_{ii}))$ is well-defined as the unique minimizer of $w_{i}^{d}+\tau_{i}$ for each $x_{ii}$ in the following proposition.
Proposition 4.1 (Optimal Radii Design).
For any $i\in S$, the optimization problem
(26) $\min_{(R_{i,c},R_{i,d})\in\mathbb{R}^{2}_{+}}w_{i}^{d}(x_{ii},R_{i,c},R_{i,d})+\tau_{i}(x_{ii},R_{i,c},R_{i,d}),$
with $w_{i}^{d}$ and $\tau_{i}$ given in (21) has a unique optimal solution $(R_{i,c}^{*},R_{i,d}^{*})$ with $R_{i,c}^{*}=R_{i,d}^{*}=R_{i}^{*}(x_{ii})$ where $R_{i}^{*}(x_{ii})$ is the unique solution to the following equation regarding $R_{i}$,
(27) $\left(\frac{b_{i}}{4v_{i}}\right)^{\frac{2}{3}}\left(\mathrm{\,erf\,}\sqrt{-\log(1-\frac{x_{ii}}{b_{i}}e^{-\frac{b_{i}-x_{ii}}{\theta_{i}}\pi R_{i}^{2}})}-\frac{2}{\sqrt{\pi}}\sqrt{-\log(1-\frac{x_{ii}}{b_{i}}e^{-\frac{b_{i}-x_{ii}}{\theta_{i}}\pi R_{i}^{2}})}(1-\frac{x_{ii}}{b_{i}}e^{-\frac{b_{i}-x_{ii}}{\theta_{i}}\pi R_{i}^{2}})\right)^{\frac{2}{3}}$
$\quad=\frac{-\log(1-\frac{x_{ii}}{b_{i}}e^{-\frac{b_{i}-x_{ii}}{\theta_{i}}\pi R_{i}^{2}})}{\pi R_{i}^{2}}$
Here, $\mathrm{\,erf\,}$ is the error function defined by $\mathrm{\,erf\,}(y)=\tfrac{2}{\sqrt{\pi}}\int_{0}^{y}e^{-s^{2}}\mathrm{d}s$.
Proof.
The optimization problem (26) is equivalent to the following problem,
(28) $\min$ $\quad\frac{u_{i}}{x_{ii}}+\frac{1}{2v_{i}}\sqrt{\frac{\theta_{i}}{b_{i}-x_{i}}}\left(\mathrm{\,erf\,}\alpha_{i}-\frac{2}{\sqrt{\pi}}\alpha_{i}e^{-\alpha_{i}^{2}}\right)+\frac{b_{i}}{2v_{i}x_{ii}\sqrt{u_{i}}}\left(\mathrm{\,erf\,}\beta_{i}-\frac{2}{\sqrt{\pi}}\beta_{i}e^{-\beta_{i}^{2}}\right)$
s.t. $\quad x_{ii}e^{-\alpha_{i}^{2}}+b_{i}e^{-\beta_{i}^{2}}=b_{i}$
over $\quad u_{i},\alpha_{i},\beta_{i}\in\mathbb{R}_{+}$
which is obtained from (26) by changes of variables as follows,
[…]
In the following theorem, we demonstrate that the sojourn time $T_{ii}(x_{ii})$ as defined in Assumption 1 is indeed a strictly increasing function of $x_{ii}$ for all $i\in\mathcal{S}$ if the proposed NN dispatch algorithm with dynamic radii is adopted.
Theorem 4.2 (Monotone Sojourn Time).
Given the optimal adjustment of matching radii $(R_{i,c}(x_{ii}),R_{i,d}(x_{ii}))$ defined by (25), i.e., $R_{i,c}(x_{ii})=R_{i,d}(x_{ii})=R_{i}^{*}(x_{ii})$, where $R_{i}^{*}(x_{ii})$ is the unique solution to (27), the following function
$f_{i}(x_{ii})=w_{i}^{d}(x_{ii},R_{i,c}(x_{ii}),R_{i,d}(x_{ii}))+\tau_{i}(x_{ii},R_{i,c}(x_{ii}),R_{i,d}(x_{ii}))$
is a strictly increasing function of $x_{ii}$ when $x_{ii}\in(0,b_{i})$ for all $i\in\mathcal{S}$. Consequently, there exists a unqiue stationary equilibrium for game $\mathcal{G}$.
Proof.
We prove the theorem using the equivalent optimization problem (28). As the objective function of (28) is smooth, it is legitimate to apply Envelop Theorem. By Envelop Theorem, one can check that the derivative of $f_{i}$ with respect to $x_{ii}$ is equal to
$\frac{\partial f_{i}}{\partial x_{ii}}$ $=\frac{\sqrt{\theta}_{i}}{4v_{i}(b_{i}-x_{ii})^{\frac{3}{2}}}\left(\mathrm{\,erf\,}\Big{(}\sqrt{\pi\frac{b_{i}-x_{ii}}{\theta_{i}}}R_{i}^{*}(x_{ii})\Big{)}-\frac{2}{\sqrt{\pi}}\sqrt{\pi\frac{b_{i}-x_{ii}}{\theta_{i}}}R_{i}^{*}(x_{ii})e^{-\frac{b_{i}-x_{ii}}{\theta_{i}}\pi(R_{i}^{*}(x_{ii}))^{2}}\right)$
$\quad+\frac{b_{i}^{\frac{2}{3}}\left(\frac{2}{\sqrt{\pi}}\beta^{*}_{i}(x_{ii})e^{-(\beta^{*}_{i}(x_{ii}))^{2}}-3\mathrm{\,erf\,}\big{(}\beta_{i}^{*}(x_{ii})\big{)}+\frac{4}{\sqrt{\pi}}\beta^{*}_{i}(x_{ii})\right)}{(4v_{i})^{\frac{2}{3}}x^{2}_{ii}\left(\mathrm{\,erf\,}\big{(}\beta_{i}^{*}(x_{ii})\big{)}-\frac{2}{\sqrt{\pi}}\beta^{*}_{i}(x_{ii})e^{-(\beta^{*}_{i}(x_{ii}))^{2}}\right)^{\frac{1}{3}}}$
where
$\beta_{i}^{*}(x_{ii})=\sqrt{-\log(1-\frac{x_{ii}}{b_{i}}e^{-\frac{b_{i}-x_{ii}}{\theta_{i}}\pi(R^{*}_{i}(x_{ii}))^{2}})},$
and $R_{i}^{*}(x_{ii})$ is the unique solution to (27). One can check that this derivative is strictly positive when $x_{ii}\in(0,b_{i})$ for all $i\in\mathcal{S}$. Then, the uniqueness of stationary equilibrium follows from Corollary 2.5. ∎
[…]
5. Simulation
Matching radius $R$ (km) Total number of arrival customers Mean waiting time for customers (min) Mean pick-up time (min) Mean waiting time for drivers (min) Mean total waiting time for customers (min) Mean total waiting time for drivers (min) Completion rate
0.5 14240 4.20 0.78 13.85 4.98 14.63 0.591
1.0 14192 2.73 1.09 7.01 3.82 8.10 0.728
1.5 14309 2.10 1.43 4.12 3.53 5.55 0.794
2.0 14501 1.72 1.72 2.68 3.44 4.40 0.826
2.5 14329 1.60 2.04 1.64 3.64 3.68 0.842
3.0 14229 1.33 2.33 1.33 3.66 3.66 0.868
3.5 14454 1.49 2.56 0.81 4.05 3.37 0.853
4.0 14275 1.33 2.84 0.67 4.17 3.51 0.864
4.5 14557 1.58 2.92 0.53 4.50 3.45 0.842
5.0 14292 1.45 3.16 0.43 4.61 3.59 0.857
Dynamic Adjustment of Radius 14355 1.27 2.40 1.21 3.67 3.61 0.870
Table 3. Comparison of Performance Metrics for Fixed Matching Radii and Dynamic Adjustment of Radius
We have theoretically proved and numerically shown that the proposed optimal adjustment of matching radius resolves the WGC problem and ensures a unique equilibrium. To further verify the effectiveness of the proposed method, we simulate the stochastic online matching process in a single region. We apply the nearest-neighbor dispatch with the optimal adjustment of matching radius where the optimal matching radius is computed for each value of driver supply rate using the equation (27). The experiments are performed on using Python 3.11 on a 16-core CPU.
The simulator is established on a space with 10 km $\times$ 10 km area. The basic setting are as follows. The simulation duration is 24 hour, i.e., 1440 min. The arrival time of customer follows a exponential distribution with rate 10 pass/min. Their locations are randomly generated. Customers are impatient and abandon the orders if the waiting time (until they are matched with the drivers) exceeds the maximum time they are willing to wait, which is exponentially distributed with mean 10 min. There are in total 200 drivers and the vehicle speed is 0.4 km/min, i.e., 24 km/h. The travel time from customer’s origin to customer’s destination is exponentially distributed with mean 20 min. After dropping a passenger, a vehicle becomes available for matching immediately and the its new location is randomly generated.
The optimal adjustment of matching radius in the simulation works as follows: we solve (27) to determine the optimal matching radius each time when a customer sends a request or when a driver becomes available where the driver supply rate is estimated by the average driver supply rate over the last hour. We compare the performance of optimal adjustment of matching radius and benchmark scenarios with fixed matching radii $R_{c}=R_{d}=R$ where the values of $R$ is taken from 0.5 km to 5 km with a step size of 0.5 km. Evaluation metrics include the mean pick-up time, mean waiting time, mean total waiting time (sum of pick-up time and waiting time) of passengers and drivers, and the completion rate (the proportion of passenger requests served successfully). The results in Table 3 show that the dynamic optimization strategy achieves a higher completion rate (0.870) than all the benchmark scenarios. The strategy also achieves a reasonably low mean total waiting time (pick up time plus waiting time) for both customers and drivers. Overall, we can claim that optimal adjustment of matching radius achieves near optimal system performance among all the simulated cases.
6. Conclusion
In this study, we have made significant strides in understanding and enhancing the operational dynamics of ride-hailing platforms through the lens of mean-field game theory. Our primary contribution lies in the introduction of a novel mean field game model, which not only systematically analyzes the intricate dynamics of these platforms but also establishes a profound connection between the equilibrium points of the game and the KKT points of an associated optimization problem. This theoretical framework offers a fresh perspective on platform operations, enabling a deeper understanding of the underlying mechanisms that govern driver and customer interactions. Furthermore, our analysis of ride-matching algorithms in the presence of customer abandonment behavior has unveiled critical insights. We have identified that these algorithms can inadvertently lead to non-monotone sojourn times, resulting in the emergence of multiple equilibria and the problematic “Wild Goose Chase” one. This WGC equilibrium, characterized by drivers engaging in inefficient pursuits of distant requests, poses a significant challenge to the overall efficiency of ride-hailing systems. To address these inefficiencies, we have proposed an innovative two-matching-radius nearest-neighbor dispatch algorithm. Our rigorous theoretical proofs and empirical validations demonstrate that the dynamic and optimal adjustment of matching radii can effectively ensure that sojourn times increase monotonically with the driver supply rate. This breakthrough not only provides a robust solution to the WGC problem but also enhances the overall efficiency and reliability of ride-hailing platforms.
We conclude this paper by envisioning a roadmap for future research that extends the boundaries of our current work. One promising avenue lies in the exploration of pricing mechanisms within ride-hailing platforms. Inspired by (Castillo et al., 2023)’s findings, which indicate that surge pricing outperforms matching adjustments in mitigating the WGC equilibrium, we aim to conduct a comprehensive and theoretical comparison using our mean field game framework. The findings in current work suggest that an optimal matching strategy can effectively eliminate WGC without reliance on any pricing designs. We expect that incorporating pricing design into our model will unveil new dimensions and insights, further enriching our understanding of ride-hailing dynamics. Another extension of our model is its application to delivery problems, where couriers, analogous to drivers in our setting, have the capability to deliver multiple parcels simultaneously. The analysis of spatial queueing networks in this context is more complex and challenging. We are particularly interested in exploring whether our approach of expressing sojourn times as functions of throughput can be adapted and applied to this more intricate setting. Such an extension could provide valuable insights into the optimization of delivery systems, contributing to enhanced efficiency and service quality.